Panama island is a hypothetical small closed economy in the northern part of Nalubia continent. Autonomous consumption in Panama dollars is 2000 and government spending = Taxes = 1000. The investment function is I=2000-25r and the money demand function is (M/P)d=Y-10r. Money supply is 1000 and the price level is 2. Let us assume that marginal propensity to consume is 0.6.

Use the information above to drag and drop answer in correct space provided.

a) IS curve equation is 16.8 and LM curve is 186.2

b) Equilibrium interest rate is 6700 and equilibrium level of national income is Y = 500 – 62.5r

c) Suppose that the government decides to double both the taxes and government spending, new equilibrium interest rate is 23.5 and new equilibrium level of national income is Y =14000 + 10r


1
Expert's answer
2021-02-15T07:54:53-0500


a) the IS equation:

"Y=\frac{1}{1-mrc}(C-mrcT)+I(r)+G)"

"16.8=\frac{1}{1-0.6}(2000-0.6\times1000+2000-25r+1000)"

"16.8=2.5(4400-25r)"

16.8=11000-62.5r

r=175.73

the IM equation:

"(\frac{M}{P})^S=(\frac{M}{P})^D=186.2"

b) r=6700

Y=500-62.5r

500-62,5r=11000-62.5r

Y=10500

C)"14000+10r=\frac{1}{1-0.6}(2000-0.6\times2000+2000-25r+2000)"

"14000+10r=2.5(3800-25r)"

14000+10r=9800-62.5r

4200=-52.5r

r=80